# Language Model Preference Evaluation with Multiple Weak Evaluators

### v3

@@ Appendix B Alternative Evaluator Configurations for Model Ranking
To address concerns about the computational cost and fairness of using 70B-level models as weak evaluators in the model ranking task, we conducted additional experiments using smaller, more comparable models. Table 5 summarizes the performance of these models, both individually and when combined using GED. The results show that GED outperforms individual evaluators even when using smaller 7B-scale models, achieving an average score of 62.70 compared to the best individual performance of 57.92 (Qwen2-7B with graph denoising). This demonstrates that the aggregation of smaller models through GED effectively enhances performance while reducing computational costs. These findings validate the versatility of GED, showing that it can provide robust and accurate rankings without relying solely on large-scale models.
@@
Table 5: Performance comparison in the model ranking task using 7B-scale models as evaluators on the AlpacaEval dataset. Higher values indicate better performance. GED achieves robust performance even with smaller evaluators.
@@
Multiple evaluator GED 57.59 61.14 61.14 67.17 66.46 62.70
Appendix C Comparison with Point-Wise Scoring Methods
Table 6: Comparison of point-wise methods and GED on response selection tasks.
Model HumanEval AlpacaEval MATH GSM8k GAIA Avg
Point-wise - Llama3-8B 60.97 28.77 74.06 82.65 10.97 51.48
Point-wise - Mistral-7B 62.83 26.93 74.02 83.21 10.22 51.44
Point-wise - Qwen2-7B 60.41 28.30 74.32 84.27 10.82 51.62
Point-wise - Majority Voting 63.72 29.42 74.71 84.36 11.32 52.70
GED (Llama3-8B, Mistral-7B, Qwen2-7B) 70.73 32.44 75.58 88.18 13.33 56.05
From the results, we observe that point-wise methods consistently underperform compared to GED across all tasks. This can be attributed to the fact that point-wise scoring assesses each response independently, without considering the relative quality of different responses. As a result, it lacks the global information available in preference-based ranking, where responses are directly compared. While majority voting improves point-wise performance slightly by aggregating scores from multiple evaluators, it remains inferior to GED, which leverages pairwise preferences to construct a more informed global ranking. Despite its limitations, the point-wise approach has an advantage in terms of computational efficiency, as it requires fewer comparisons than pairwise methods. This trade-off suggests that the choice between point-wise and preference-based scoring should depend on the specific application: point-wise scoring may be preferable in large-scale settings where efficiency is critical, while GED is more effective for achieving higher-quality rankings.
Appendix D Impact of Evaluator Quantity on Denoising Quality
We investigated how the number of evaluators affects the denoising quality in GED by conducting experiments using different numbers of evaluators in the response selection setting. Specifically, we evaluated the performance of GED when aggregating preferences from two, three, and four evaluators. The evaluators used were Llama3-8B, Mistral-7B, Qwen2-7B, and Gemma-9B [^3] https://huggingface.co/google/gemma-2-9b-it.
Table 7: Performance comparison of GED with varying numbers of evaluators across five benchmarks. Increasing the number of evaluators enhances denoising quality, reflected in improved performance metrics.
Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg.
Llama3-8B, Mistral-7B 69.21 31.87 74.97 86.92 12.51 55.10
Llama3-8B, Mistral-7B, Qwen2-7B 70.73 32.44 75.58 88.18 13.33 56.05
As shown in Table 7, increasing the number of evaluators consistently improves the denoising quality of GED, as reflected by higher performance across all benchmarks. For instance, the average performance improves from 55.10% with two evaluators to 56.39% with four evaluators. This enhancement can be attributed to the diverse perspectives and complementary strengths of multiple evaluators, contributing to a more robust and accurate aggregation of preferences. These results highlight the importance of selecting a diverse and capable set of evaluators to enhance GED’s effectiveness. Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
@@ Appendix D Impact of Evaluator Quantity on Denoising Quality
Appendix E Construction of the Preference Graph
In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 1 and 2. These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
Appendix F The Importance of Addressing Cyclic Inconsistency
Cyclic inconsistencies in preference data introduce contradictions that undermine meaningful ranking and evaluation. Many applications, including alignment optimization (Ouyang et al., 2022; Song et al., 2024), recommendation systems (Hou et al., 2024), and model evaluation (Liu et al., 2024), rely on transitive preferences for stability and interpretability. For example, OpenAI’s RLHF pipelines (Ouyang et al., 2022) construct training data based on transitive preferences (e.g., A $\succ$ B $\succ$ C), which ensures consistency in model learning. However, if preference cycles exist (e.g., A $\succ$ B $\succ$ C $\succ$ A), the resulting contradictions lead to instability and degrade training performance. In our experiments, removing cycles and transforming preference graphs into directed acyclic graphs (DAGs) improved results across multiple tasks, including response selection, model ranking, and instruction tuning. These improvements were consistent across widely used benchmarks, suggesting that most observed cycles are artifacts of noise rather than meaningful violations of transitivity. Ensuring acyclicity enhances dataset reliability and improves model effectiveness in ranking and evaluation tasks.
Appendix G Proof of Theorem 1
@@
Proof.
For brevity, we consider all edges in $G$ have weights equal to 1 and all weights in $\widehat{G}$ are divided by $N$. By construction, we can notice that for each $(u\rightarrow v)\in A$, the weight $w_{\widehat{G}}(v\rightarrow u)$ can be viewed as an empirical estimate of $\delta_{1}$. Then, we claim that the following two events can imply $G\subseteq\mathrm{MAS}(\widehat{G})$:
@@
To see this, first by Lemma 1, we know that for any pair of nodes $(u,v)$, $\mathrm{MAS}(\widehat{G})$ will contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$. [^4] There is non-zero probability that some edges in $\widehat{G}$ will have zero weight, but we treat them as existing for the ease of argument. That is, we allow only $\widehat{G}$ to contain zero-weight edges. Therefore, for any $(u\rightarrow v)\in A$, $\mathrm{MAS}(\widehat{G})$ will contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$. Then, since $\mathcal{E}_{1}$ holds, $\delta_{1}=0.5-\epsilon<0.5$ and $w_{\widehat{G}}(u\rightarrow v)+w_{\widehat{G}}(v\rightarrow u)=1$, we have $w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\geq\epsilon$. Furthermore, since $\mathcal{E}_{2}$ holds, for $(u,v)$ such that $(u\rightarrow v),(v\rightarrow u)\notin A$, arbitrary way of edge removing among these nodes can influence the total edge weights by at most $\epsilon$. Therefore, when applying the denoising operation to $\widehat{G}$, for any $(u\rightarrow v)\in A$, only $(u\rightarrow v)$ will be kept in $\mathrm{MAS}(\widehat{G})$, which makes $G\subseteq\mathrm{MAS}(\widehat{G})$. As a result, we have $\mathbb{P}(\mathcal{E}_{1}\cap\mathcal{E}_{2})\leq\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)$.
@@
$\mathbb{P}\left(\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}\right)\geq$ $1-2\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right)$
$\implies\mathbb{P}(\mathcal{E}_{2})\geq$ $1-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
@@
$\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)\geq$ $\mathbb{P}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right).$
Lemma 1.
@@
We will then use contradiction to show it is impossible for $\mathrm{MAS}(G)$ to contain neither $(u\rightarrow v)$ nor $(v\rightarrow u)$. Suppose this is true instead. Then, since $\mathrm{MAS}(G)$ is a maximum acyclic subgraph, adding either $(u\rightarrow v)$ or $(v\rightarrow u)$ to $\mathrm{MAS}(G)$ will make it cyclic. That is, $\mathrm{MAS}(G)$ contains a path that goes from $v$ to $u$; meanwhile, it also contains a path from $u$ to $v$. As a result, it contains a cycle that goes from $v$ to $u$ and then from $u$ to $v$, which contradicts with the fact that $\mathrm{MAS}(G)$ is a maximum acyclic subgraph. Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$. ∎
Appendix H Impact of Evaluator Weighting on GED Performance
@@
In our theoretical analysis, GED assumes equal weighting of edges in the preference graphs. However, in practical scenarios, evaluators may have varying levels of reliability or expertise. Incorporating evaluator-specific confidence scores or performance metrics could enhance the effectiveness of GED. To investigate this, we conducted experiments using a weighted version of GED, referred to as WeightGED, under the Response Selection setting. In WeightGED, we assigned weights to evaluators based on their individual performance on specific datasets. For example, on the GSM8k dataset, the response selection accuracies for Llama3-8B, Mistral-7B, and Qwen2-7B were 62.19%, 67.24%, and 61.58%, respectively. These accuracies were normalized to compute evaluator weights:
@@
[…]
These weights were used to scale the contributions of each evaluator’s preferences during graph construction. We compared the performance of GED and WeightGED across multiple benchmarks, as presented in Table 8.
Table 8: Performance comparison of GED and WeightGED using 7B-scale evaluators across five benchmarks. WeightGED marginally outperforms GED, demonstrating the potential benefits of incorporating evaluator-specific weights.
Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
GED 70.73 32.44 75.58 88.18 13.33 56.05
7B Evaluators WeightGED 70.97 32.67 75.71 88.24 13.56 56.23
GED 73.21 59.87 82.49 86.43 16.27 63.65
GPT Evaluators WeightGED 74.52 61.71 83.93 87.84 17.32 65.06
As shown in Table 8, WeightGED achieves marginal but consistent improvements over GED across all benchmarks. For instance, using 7B-scale evaluators, WeightGED improves the average performance from 56.05% to 56.23%. Similarly, with GPT evaluators, the average performance increases from 63.65% to 65.06%. These results suggest that incorporating evaluator-specific weights based on performance metrics can enhance the effectiveness of GED. Furthermore, we conducted additional experiments using stronger evaluators such as GPT-3.5, GPT-4-o-mini, and GPT-4-o. The weights were computed based on their respective performance accuracies, following the same normalization procedure. The improvements observed with these evaluators reinforce the potential of weighting schemes in enhancing GED. In summary, our preliminary findings indicate that integrating evaluator-specific confidence scores or performance metrics is a promising direction for future work. Systematically designing and optimizing these weighting schemes could lead to more robust and accurate evaluation frameworks.
Appendix I Expanded Related Work on Preference Denoising for LLMs
This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs. We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing. Below, we provide a detailed comparison of GED with representative methods from these approaches. Within-model approaches, such as robust DPO (rDPO) (Chowdhury et al., 2024) and conservative DPO (cDPO) (Mitchell, 2023), focus on integrating denoising mechanisms directly within the preference modeling process. These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment. While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility. In contrast, GED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models. Modular approaches like CURATRON (Naresh et al., 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion. While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of GED. By leveraging a graph-based framework, GED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use. Additionally, GED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions. Furthermore, the ensemble mechanism in GED demonstrates the novel insight that combining small evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
Appendix J Impact of Evaluator Selection on GED Performance

### v4

@@ Appendix E Alternative Evaluator Configurations for Model Ranking
To address concerns about the computational cost and fairness of using 70B-level models as preference evaluators in the model ranking task, we conducted additional experiments using smaller, more comparable models. Table 5 summarizes the performance of these models, both individually and when combined using PGED. The results show that PGEDoutperforms individual evaluators even when using smaller 7B-scale models, achieving an average score of 62.70 compared to the best individual performance of 57.92 (Qwen2-7B with graph denoising). This demonstrates that the aggregation of smaller models through PGEDeffectively enhances performance while reducing computational costs. These findings validate the versatility of PGED, showing that it can provide robust and accurate rankings without relying solely on large-scale models.
@@
Table 5: Performance comparison in the model ranking task using 7B-scale models as evaluators on the AlpacaEval dataset. Higher values indicate better performance. PGEDachieves robust performance even with smaller evaluators.
@@
Multiple evaluator PGED 57.59 61.14 61.14 67.17 66.46 62.70
@@
Appendix F Comparison with Point-Wise Scoring Methods
Table 6: Comparison of point-wise methods and PGEDon response selection tasks.
Model HumanEval AlpacaEval MATH GSM8k GAIA Avg
Point-wise - Llama3-8B 60.97 28.77 74.06 82.65 10.97 51.48
Point-wise - Mistral-7B 62.83 26.93 74.02 83.21 10.22 51.44
Point-wise - Qwen2-7B 60.41 28.30 74.32 84.27 10.82 51.62
Point-wise - Majority Voting 63.72 29.42 74.71 84.36 11.32 52.70
PGED(Llama3-8B, Mistral-7B, Qwen2-7B) 70.73 32.44 75.58 88.18 13.33 56.05
From the results, we observe that point-wise methods consistently underperform compared to PGEDacross all tasks. This can be attributed to the fact that point-wise scoring assesses each response independently, without considering the relative quality of different responses. As a result, it lacks the global information available in preference-based ranking, where responses are directly compared. While majority voting improves point-wise performance slightly by aggregating scores from multiple evaluators, it remains inferior to PGED, which leverages pairwise preferences to construct a more informed global ranking. Despite its limitations, the point-wise approach has an advantage in terms of computational efficiency, as it requires fewer comparisons than pairwise methods. This trade-off suggests that the choice between point-wise and preference-based scoring should depend on the specific application: point-wise scoring may be preferable in large-scale settings where efficiency is critical, while PGEDis more effective for achieving higher-quality rankings.
Appendix G Impact of Evaluator Quantity on Denoising Quality
We investigated how the number of evaluators affects the denoising quality in PGEDby conducting experiments using different numbers of evaluators in the response selection setting. Specifically, we evaluated the performance of PGEDwhen aggregating preferences from two, three, and four evaluators. The evaluators used were Llama3-8B, Mistral-7B, Qwen2-7B, and Gemma-9B.
Table 7: Performance comparison of PGEDwith varying numbers of evaluators across five benchmarks. Increasing the number of evaluators enhances denoising quality, reflected in improved performance metrics.
Evaluators Set HumanEval AlpacaEval MATH GSM8k GAIA Avg.
Llama3-8B, Mistral-7B 69.21 31.87 74.97 86.92 12.51 55.10
Llama3-8B, Mistral-7B, Qwen2-7B 70.73 32.44 75.58 88.18 13.33 56.05
As shown in Table 7, increasing the number of evaluators consistently improves the denoising quality of PGED, as reflected by higher performance across all benchmarks. For instance, the average performance improves from 55.10% with two evaluators to 56.39% with four evaluators. This enhancement can be attributed to the diverse perspectives and complementary strengths of multiple evaluators, contributing to a more robust and accurate aggregation of preferences. These results highlight the importance of selecting a diverse and capable set of evaluators to enhance PGED’s effectiveness. Incorporating more evaluators allows the denoising process to better identify and mitigate inconsistencies in the preference graphs, leading to improved overall performance in response selection tasks.
Appendix H Construction of the Preference Graph
In this section, we provide a detailed explanation of the process used to construct the preference graph sets for both the model ranking and response selection tasks, as outlined in Algorithms 2 and 3. These algorithms form the backbone of our method, enabling the representation of pairwise preferences as directed graphs, which are essential for downstream aggregation and ranking.
@@ Appendix H Construction of the Preference Graph
Appendix I The Importance of Addressing Cyclic Inconsistency
Cyclic inconsistencies in preference data introduce contradictions that undermine meaningful ranking and evaluation. Many applications, including alignment optimization (Ouyang et al. 2022; Song et al. 2024), recommendation systems (Hou et al. 2024), and model evaluation (Liu et al. 2024), rely on transitive preferences for stability and interpretability. For example, OpenAI’s RLHF pipelines (Ouyang et al. 2022) construct training data based on transitive preferences (e.g., A $\succ$ B $\succ$ C), which ensures consistency in model learning. However, if preference cycles exist (e.g., A $\succ$ B $\succ$ C $\succ$ A), the resulting contradictions lead to instability and degrade training performance. In our experiments, removing cycles and transforming preference graphs into directed acyclic graphs (DAGs) improved results across multiple tasks, including response selection, model ranking, and instruction tuning. These improvements were consistent across widely used benchmarks, suggesting that most observed cycles are artifacts of noise rather than meaningful violations of transitivity. Ensuring acyclicity enhances dataset reliability and improves model effectiveness in ranking and evaluation tasks.
Appendix J Proof of Theorem 1
@@ Appendix I The Importance of Addressing Cyclic Inconsistency
Proof J.1
[…]
@@
[…]
@@
$\mathbb{P}\left(\left|w_{\widehat{G}}(u\rightarrow v)-w_{\widehat{G}}(v\rightarrow u)\right|\leq\frac{\epsilon}{U}\right)$
$\geq 1-2\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right)$
$\implies\mathbb{P}(\mathcal{E}_{2})$ $\geq 1-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right),$
@@
$\mathbb{P}\left(G\subseteq\mathrm{MAS}(\widehat{G})\right)\geq\mathbb{P}\left(\mathcal{E}_{1}\cap\mathcal{E}_{2}\right)$
$\geq 1-2|A|\exp\left(-\frac{N\epsilon^{2}}{2}\right)-2U\exp\left(-\frac{N\epsilon^{2}}{6U^{2}\delta_{2}+2U\epsilon}\right).$
Lemma 1
@@
We will then use contradiction to show it is impossible for $\mathrm{MAS}(G)$ to contain neither $(u\rightarrow v)$ nor $(v\rightarrow u)$. Suppose this is true instead. Then, since $\mathrm{MAS}(G)$ is a maximum acyclic subgraph, adding either $(u\rightarrow v)$ or $(v\rightarrow u)$ to $\mathrm{MAS}(G)$ will make it cyclic. That is, $\mathrm{MAS}(G)$ contains a path that goes from $v$ to $u$; meanwhile, it also contains a path from $u$ to $v$. As a result, it contains a cycle that goes from $v$ to $u$ and then from $u$ to $v$, which contradicts with the fact that $\mathrm{MAS}(G)$ is a maximum acyclic subgraph. Therefore, $\mathrm{MAS}(G)$ must contain exactly one of $(u\rightarrow v)$ and $(v\rightarrow u)$.
Appendix K Impact of Evaluator Weighting on PGED
@@
In our theoretical analysis, PGEDassumes equal weighting of edges in the preference graphs. However, in practical scenarios, evaluators may have varying levels of reliability or expertise. Incorporating evaluator-specific confidence scores or performance metrics could enhance the effectiveness of PGED. To investigate this, we conducted experiments using a weighted version of PGED, referred to as WeightGED, under the Response Selection setting. In WeightGED, we assigned weights to evaluators based on their individual performance on specific datasets. For example, on the GSM8k dataset, the response selection accuracies for Llama3-8B, Mistral-7B, and Qwen2-7B were 62.19%, 67.24%, and 61.58%, respectively. These accuracies were normalized to compute evaluator weights:
$\text{weight}(\text{Llama3-8B})$ $=\frac{62.19}{62.19+67.24+61.58}=0.326,$
$\text{weight}(\text{Mistral-7B})$ $=\frac{67.24}{62.19+67.24+61.58}=0.352,$
$\text{weight}(\text{Qwen2-7B})$ $=\frac{61.58}{62.19+67.24+61.58}=0.322.$
These weights were used to scale the contributions of each evaluator’s preferences during graph construction. We compared the performance of PGEDand WeightGED across multiple benchmarks, as presented in Table 8.
Table 8: Performance comparison of PGEDand WeightGED using 7B-scale evaluators across five benchmarks. WeightGED marginally outperforms PGED, demonstrating the potential benefits of incorporating evaluator-specific weights.
Method HumanEval AlpacaEval MATH GSM8k GAIA Avg.
PGED 70.73 32.44 75.58 88.18 13.33 56.05
7B Evaluators WeightGED 70.97 32.67 75.71 88.24 13.56 56.23
PGED 73.21 59.87 82.49 86.43 16.27 63.65
GPT Evaluators WeightGED 74.52 61.71 83.93 87.84 17.32 65.06
As shown in Table 8, WeightGED achieves marginal but consistent improvements over PGEDacross all benchmarks. For instance, using 7B-scale evaluators, WeightGED improves the average performance from 56.05% to 56.23%. Similarly, with GPT evaluators, the average performance increases from 63.65% to 65.06%. These results suggest that incorporating evaluator-specific weights based on performance metrics can enhance the effectiveness of PGED. Furthermore, we conducted additional experiments using stronger evaluators such as GPT-3.5, GPT-4-o-mini, and GPT-4-o. The weights were computed based on their respective performance accuracies, following the same normalization procedure. The improvements observed with these evaluators reinforce the potential of weighting schemes in enhancing PGED. In summary, our preliminary findings indicate that integrating evaluator-specific confidence scores or performance metrics is a promising direction for future work. Systematically designing and optimizing these weighting schemes could lead to more robust and accurate evaluation frameworks.
Appendix L Expanded Related Work on Preference Denoising for LLMs
This work situates itself within the broader field of denoising preference evaluations for LLMs, addressing inconsistencies and noise in preference graphs. We acknowledge that existing literature has explored two primary approaches to this challenge: within-model preference modeling and modular pre-processing. Below, we provide a detailed comparison of PGED with representative methods from these approaches. Within-model approaches, such as robust DPO (rDPO) (Chowdhury, Kini, and Natarajan 2024) and conservative DPO (cDPO) (Mitchell 2023), focus on integrating denoising mechanisms directly within the preference modeling process. These methods incorporate regularization techniques to mitigate the effects of noisy or adversarial preference data during model alignment. While these approaches are effective in refining the preference modeling pipeline, they are tightly coupled with specific models and tasks, limiting their versatility. In contrast, PGED is a modular framework that operates as a pre-processing step, denoising preference graphs before downstream tasks, making it adaptable to a wide range of applications and models. Modular approaches like CURATRON (Naresh, Tulabandhula et al. 2024) address noise and missing comparisons in preference datasets using techniques such as low-rank matrix completion. While CURATRON effectively mitigates certain types of noise, it does not explicitly target cyclic inconsistencies (e.g., A $\succ$ B, B $\succ$ C, C $\succ$ A), which are a critical focus of PGED. By leveraging a graph-based framework, PGED detects and removes such cycles through its denoising process, ensuring that the resulting preference graph is acyclic and thus more consistent and reliable for downstream use. Additionally, PGED distinguishes itself by providing theoretical guarantees for recovering the ground truth preference DAG under certain conditions. Furthermore, the ensemble mechanism in PGED demonstrates the novel insight that combining small evaluators can surpass the performance of a single stronger evaluator, a feature not emphasized in the aforementioned methods.
Appendix M Theoretical Analysis of Graph Denoising
Let $G=(V,A,w)$ be the aggregated digraph, with $w:A\to\mathbb{R}_{\geq 0}$. Write $m\triangleq|\{(u,v)\in A:w(u,v)>0\}|$. For $u\in V$, define
$d^{+}(u)$ $\triangleq\sum_{(u,v)\in A}w(u,v),d^{-}(u)\triangleq\sum_{(v,u)\in A}w(v,u),$
$d(u)$ $\triangleq d^{+}(u)+d^{-}(u),\delta(u)\triangleq d^{+}(u)-d^{-}(u).$
Let $s=s_{1}\|s_{2}$ be the order returned by Algorithm 1, and
$R(s)$ $\triangleq\{(v_{j}\to v_{i})\in A:\ j>i\text{ in }s\},$
$G^{\prime}$ $\triangleq(V,A\setminus R(s),w).$
Proposition 1 (Acyclicity)
$G^{\prime}$ is a directed acyclic graph.
Proof M.1
We prove by contradiction. Assume $G^{\prime}$ contains a directed cycle $C=(u_{1}\!\to u_{2}\!\to\cdots\!\to u_{\ell}\!\to u_{1})$. Since $G^{\prime}$ keeps exactly the forward arcs w.r.t. the total order $s$, every arc $(u_{t}\!\to u_{t+1})$ in $C$ must satisfy $\mathrm{pos}_{s}(u_{t})<\mathrm{pos}_{s}(u_{t+1})$, where $\mathrm{pos}_{s}(\cdot)$ is the index in $s$. Summing these strict inequalities around the cycle yields
$\mathrm{pos}_{s}(u_{1})<\mathrm{pos}_{s}(u_{2})<\cdots<\mathrm{pos}_{s}(u_{\ell})<\mathrm{pos}_{s}(u_{1}),$
which is impossible. Therefore $G^{\prime}$ has no directed cycle.
Assumption 1 (Bounded integral weights)
There exists a fixed integer $k\geq 1$ (the number of evaluators) such that $w(u,v)\in\{0,1,\dots,k\}$ for all $(u,v)\in A$.
Assumption 1 holds in our ensemble setting because each evaluator contributes at most one unit to any arc weight.
Data structures.
We maintain: (i) a queue $Q^{-}$ of sources ($d^{-}(u)=0$); (ii) a queue $Q^{+}$ of sinks ($d^{+}(u)=0$); (iii) $\delta$-buckets $\{B_{\Delta}\}_{\Delta=-M}^{M}$, where $M\triangleq\max_{u\in V}d(u)$. Each bucket is a doubly linked list of vertices with current $\delta(u)=\Delta$ and $d^{\pm}(u)>0$. We also keep an integer pointer $p$ storing the maximum index $\Delta$ with $B_{\Delta}\neq\emptyset$. All vertices are in exactly one of $^{-}, Q+Q^{+}$, or some $B_{\Delta}$.
Lemma 2 (Range and update magnitude)
Under Assumption 1: (a) $M=\mathcal{O}(m)$; (b) when removing a vertex $u$, for any neighbor $v$ we have
$\delta_{\text{new}}(v)$ $=\delta_{\text{old}}(v)+w(u,v)-w(v,u),$
$\Rightarrow\quad|\delta_{\text{new}}(v)-\delta_{\text{old}}(v)|$ $\leq k.$
Proof M.2
(a) Since each incident arc contributes at most $k$ to $d^{\pm}(\cdot)$,
$d(u)=\sum_{(u,x)}w(u,x)+\sum_{(x,u)}w(x,u)\;\leq\;k\cdot\deg(u),$
hence $M=\max_{u}d(u)\leq k\cdot\max_{u}\deg(u)\leq k\cdot m=\mathcal{O}(m)$. (b) When $u$ is removed, exactly the arcs $(u,v)$ and $(v,u)$ (if present) are deleted among $v$’s incidences. Thus $d^{-}(v)$ decreases by $w(u,v)$, $d^{+}(v)$ decreases by $w(v,u)$, and $\delta(v)=d^{+}(v)-d^{-}(v)$ increases by $w(u,v)$ and decreases by $w(v,u)$, as claimed. By Assumption 1, $w(u,v),w(v,u)\leq k$, hence the $|\cdot|$ bound.
Lemma 3 (Bucket maintenance in $\mathcal{O}(1)$ amortized time)
Under Assumption 1 , the following hold.
  1. 1.
    Initialization. Computing $d^{\pm}(\cdot)$ and placing vertices into $^{-}, Q+Q^{+}$, or $B_{\Delta}$ takes $\mathcal{O}(m)$ time and $\mathcal{O}(m)$ space.
  2. 2.
    Updates. Removing a vertex $u$ and processing all its incident arcs take time $\mathcal{O}(1+\deg(u))$: every neighbor $v$ is updated once, and $v$ moves from $B_{\Delta}$ to $B_{\Delta^{\prime}}$ with $|\Delta^{\prime}-\Delta|\leq k$; enqueuing/dequeuing from $Q^{\pm}$ or $B_{\Delta}$ is $\mathcal{O}(1)$ via doubly linked lists.
  3. 3.
    Selecting $\max\delta$. Maintaining pointer $p$ to the largest index with $B_{p}\neq\emptyset$ costs $\mathcal{O}(1)$ amortized per removal.
Proof M.3
(1) We make one pass over adjacency lists: for $(x,y)$ add $w(x,y)$ to $d^{+}(x)$ and to $d^{-}(y)$. This is $\mathcal{O}(m)$. Classification into $Q^{\pm}$ or some $B_{\Delta}$ is then $\mathcal{O}(n)$, dominated by $\mathcal{O}(m)$ since a connected digraph has $m\geq n-1$. (2) When removing $u$, we unlink $u$ from its current container in $\mathcal{O}(1)$. For each incident arc touching $u$, we visit the unique neighbor $v$ once, adjust $d^{\pm}(v)$ by Lemma 2, recompute $\delta(v)$, and move $v$:
@@
Each move is pointer manipulation in a doubly linked list, costing $\mathcal{O}(1)$. Hence the work per removal is $\mathcal{O}(1+\deg(u))$ and each arc is processed exactly once (when the first endpoint is removed). (3) Amortized bound for $p$. Let $p_{t}$ be the pointer after the $t$-th removal. There are two kinds of motion:
Downward motion. Whenever $B_{p_{t}}$ becomes empty, we decrement $p_{t}$ until finding the next nonempty bucket. Since $p$ never moves below M-M and never increases during this downward scan, the total number of decrements over the entire execution is at most $2M=\mathcal{O}(m)$.
Upward motion. Deleting $u$ can increase some $\delta(v)$, but by Lemma 2 the new maximum bucket index can exceed the previous maximum by at most $k$:
$\max_{x\in V}\delta_{t}(x)$ $\leq\max\Bigl\{\max_{x\in V}\delta_{t-1}(x),\ \max_{v\in N(u)}\delta_{t-1}(v)+k\Bigr\}$
$\leq p_{t-1}+k.$
We therefore scan upward by at most $k$ to reestablish $p_{t}$. Since $k$ is a fixed constant and there are $|V|$ removals, the total upward scans cost $\mathcal{O}(k|V|)=\mathcal{O}(|V|)$. For connected digraphs, $|V|\leq m+1$, hence $\mathcal{O}(|V|)=\mathcal{O}(m)$. Combining both motions gives $\mathcal{O}(1)$ amortized per removal.
Theorem 2 (Linear time and space)
Under Assumption 1, Algorithm 1 runs in $\mathcal{O}(m)$ time and uses $\mathcal{O}(m)$ space.
Proof M.4
Time. Initialization is $\mathcal{O}(m)$ by Lemma 3(1). Across the execution, each vertex is removed once. By Lemma 3(2), the cost of removing $u$ is $\mathcal{O}(1+\deg(u))$. Summing over all $u$ and noting $\sum_{u}\deg(u)=2m$, we obtain
$\sum_{u\in V}\mathcal{O}(1+\deg(u))$ $=\mathcal{O}\bigl(|V|+\sum_{u}\deg(u)\bigr)$
$=\mathcal{O}(|V|+m)$
$=\mathcal{O}(m),$
since a connected digraph has $|V|\leq m+1$. The amortized $\mathcal{O}(1)$ selection of the maximum-$\delta$ bucket follows from Lemma 3(3), and is already charged in the removal step. Therefore the overall running time is $\mathcal{O}(m)$.
Space. Adjacency lists and degree/weight arrays take $\mathcal{O}(m)$. The queues $Q^{\pm}$ store disjoint subsets of $V$; the buckets $\{B_{\Delta}\}$ occupy an array of size $2M+1=\mathcal{O}(m)$ with total stored vertices $|V|$. All pointers and auxiliary counters are $\mathcal{O}(1)$ each. Hence the working memory is $\mathcal{O}(m)$.
Remark 1 (Arbitrary real weights)
If $w$ is real-valued without a constant bound on per-arc increments, the bucket update in Lemma 3 no longer yields $\mathcal{O}(1)$. Replacing buckets by a binary max-heap over current $\delta(u)$ preserves correctness and yields $\mathcal{O}(m\log n)$ time and $\mathcal{O}(m)$ space.
Appendix N Impact of Evaluator Selection on PGED